OpenStudy (zackattack350):

for the statement find the constant of variation and the variation equation. y varies jointly as x and the cube of z;y=24 when x=1 and z=2

OpenStudy (zackattack350):

@lochana

OpenStudy (lochana):

hmm.. okay. like they say in the question, y is given by sum of x , z^3 and some constant(say c). we get y = x + z^3 + c y= 24; x=1; z=2 24 = 1 + 2^3 + c c = 24 - 9 c = 15 constant = 15.

OpenStudy (zackattack350):

so if thats the constant using the equation at the top... how would i answer what the variation equation is? i have in the assignment y=(blank)xz^3

OpenStudy (lochana):

this is the best way that goes with your question. one can argue in many ways saying this is not to be true. but you can't find the constant if you go more deeper than my way. for example; one can ask , how do you exactly say coefficients of x , z is 1. that is not given. but I assume that to be 1. So this problem is not quite good one.:)

OpenStudy (zackattack350):

1 is a good assumption in my opinion. But i still have to answer what the equation is

OpenStudy (lochana):

but wait

OpenStudy (lochana):

I didn't get "jointly" here. After googling, I understood that jointly means multiplication.:) (english is not my native language) okay so y = cxz^3

OpenStudy (zackattack350):

oh! thank you i was confused about that as well.. Thanks for your help! ill tag you if i need you again :)

OpenStudy (lochana):

great